Martha s preference over two goods x and y is represented by the utility function U = 5xy. Her budget constraint is given by 5x +y =30. Plot budget constraint with y on the vertical - axis on the horizontal axis.. Find the marginal utility of x and y...Find the MRS.. Find the MRT...Given her budget how much of each good should she consume in order to maximize her utility show your work.


1
Expert's answer
2022-03-10T10:14:00-0500

a)

b)i) "Mu_x= 5Y"

ii) "Mu_y= 5X"

c) Maximum combination for utility maximization

"\frac{Mu_x}{Mu_y}= \frac{P_x}{P_y}"


"\frac{5Y}{5X}= \frac{P_x}{P_y}"


"\frac{Y}{X}= \frac{P_x}{P_y}"


"X= \frac{P_yY}{P_x}" .........(i)


"Y= \frac{P_xX}{P_y}" -...........(ii)

Plug into the budget constraint

5X+Y=30

5("\frac{P_yY}{P_x})+Y=30"


"\frac{5P_yY}{P_x}+Y=30"

"5P_yY+P_xY=30P_x"

"Y(5P_y+P_x)=30P_x"

"Y^*= \frac{30P_x}{5P_y+P_x}"

"5X+ \frac{P_xX}{P_y}=30"

"5P_yX+P_xX=30P_y"

"X(5P_y+P_x)= 30P_y"

"X^*= \frac{30P_y}{5P_y+P_x}"